OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4)

OpenStudy (anonymous):

3x+y=20

OpenStudy (anonymous):

I put it in slope-intercept form first y-2=-3(x-6)=y=-3x+20 Then i put it in standard form 3x+y=20

OpenStudy (anonymous):

brilliant! new to openstudy, medal?

OpenStudy (anonymous):

Sure just click the best response button

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

take it easy buddy!

OpenStudy (anonymous):

Ight bruh!

OpenStudy (anonymous):

help with slope intercept form? B(0, 0), C(4, -4)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y-(-4)=-1(x-4) y+4=-x+4 y=-x

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

-4x+y=-1?

OpenStudy (anonymous):

no y=-x

OpenStudy (anonymous):

what are the answer choices

OpenStudy (anonymous):

slope-intercept=y=mx+b

OpenStudy (anonymous):

ok ok, C(0, 4), m = 0?

OpenStudy (anonymous):

m=-1 get the slop by the formula y2-y1/x2-x1

OpenStudy (anonymous):

so it would be -4-0/4-0=-4/4=-1

OpenStudy (anonymous):

what formula for intercept? or how do i find it

OpenStudy (anonymous):

start with point-slop: y-y1=m(x-x1)